OpenStudy (anonymous):

a particle moves along the x-axis so that its velocity at t >/= 0 is given by v(t)=1-sin(2pi(t)) - find acceleration a(t) of the particle at any time t

OpenStudy (amistre64):

the derivative of velocity is acceleration

OpenStudy (amistre64):

you got a teacher telling you to post this here by chance? this is coming up alot

OpenStudy (anonymous):

yeh,, it's for school

OpenStudy (anonymous):

how do i find the derivative?

OpenStudy (amistre64):

the derivative is explained in your material :) \[\frac{d}{dt}(1-sin(2\pi\ t))= \frac{d}{dt}1-\frac{d}{dt}sin(2\pi\ t)\]

OpenStudy (amistre64):

the derivative of a constant is?

OpenStudy (anonymous):

so it would be -cos 2 pi?

OpenStudy (amistre64):

yes, except for one minor detail; the "t" parts

OpenStudy (amistre64):

sin(nt) pops out the derivative of nt

OpenStudy (amistre64):

chain rule

OpenStudy (anonymous):

sorry i dont understand

OpenStudy (anonymous):

how do you find the derivative of t/

OpenStudy (amistre64):

\[-\frac{d}{dt}sin(2\pi\ t)=-cos(2\pi\ t)*\frac{d}{dt}2\pi\ t\]

OpenStudy (anonymous):

1-cos2pi sin t?

OpenStudy (anonymous):

o ok! got it

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

so final answer is 1-cos2pi x 2pi(t)

OpenStudy (amistre64):

2pi t is just a constant * t; kt the derivative of say: kt is k sooooo, 2pi t derives to 2pi

OpenStudy (anonymous):

so 1-cos2pi x 2pi

OpenStudy (amistre64):

and the "1" part derived to zero so its not included in the end \[\frac{d}{dt}(1-sin(2\pi\ t))=-2\pi\ cos(2\pi\ t) \]

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

the follow up question is find all values of t 0</= t </= 2 for which the particle is at rest

OpenStudy (anonymous):

do i set the original equation = 0

OpenStudy (amistre64):

when the derivative = 0; we are at rest; the derivative of position is velocity; and 0 velocity is resting

OpenStudy (amistre64):

in this case; when cos(2pi t) = 0 we are at rest

OpenStudy (anonymous):

so we set the original to 0 to find at rest/

OpenStudy (amistre64):

2pi t = pi/2 t = 1/4 2pi t = 3pi/2 t = 3/4

OpenStudy (amistre64):

no, the original is postiion. set the derivative to zero to determine when there is no velocity; when its standing still; when its not moving; when it is at rest

OpenStudy (amistre64):

position tells us WHERE we are derivative of position = velocity; tells us HOW FAST we are moving

OpenStudy (anonymous):

but isnt our derivative 2pi cost (2pi (t)

OpenStudy (amistre64):

yes; and when is that equal to 0?

OpenStudy (amistre64):

cos = 0 at pi/2 (90) and 3pi/2 (270)

OpenStudy (amistre64):

so when 2pit = pi/2 OR 2pit = 3pi/2 we are at 0 velocity

OpenStudy (amistre64):

might wanna check to make sure we hit all the 2 values between 0 and 2 tho; theres prolly a few more to hit

OpenStudy (amistre64):

all the "t" values between .... that is

OpenStudy (amistre64):

2pi t = (2n+1)pi/2 2 t = (2n+1)/2 t = (2n+1)/4 0 < (2n+1)/4 < 2 0 < 2n+1 < 8 -1 < 2n < 7 -1/2 < n < 7/2

OpenStudy (anonymous):

its 0 to 2 thanks